OpenStudy (wannabegurl):

: what is the reciprocal of 13

OpenStudy (nurali):

The reciprocal of any term, a = 1/a The reciprocal of x/y = 1/(x/y) = y/x The reciprocal of 13 = a/13

OpenStudy (nurali):

if correct please click best response.